# Diffusiophoretic corner flows

### v1

@@ Abstract
We study flows induced in a two-dimensional corner by the chemical activity of the confining boundaries. A catalytic reaction on the surfaces drives the diffusiophoretic flow of viscous fluid in the whole domain. We show that phoretically active sectors can generate eddies in the corner, similar to Moffatt eddies, induced mechanically in a similar confinement. We analyse the exact solution of the diffusion problem in the wedge geometry, coupled to diffusiophoretic flow generation in the slip-velocity formulation, to ultimately arrive at the stream function of the flow. We discuss cases in which simple exact solutions for the flow are available. Our work can inspire the design of microscale mixing devices involving dead-end pores in microfluidic devices and can serve as benchmarks for numerical solutions for complex geometries in (diffusio)phoretic systems.
@@ I Introduction
A promising approach exploits phoretic mechanisms, in which surface-generated gradients (of concentration, temperature, etc.) induce effective slip flow on confining surfaces, that in turn gives rise to a bulk flow [4]. Diffusiophoresis, the motion of particles and fluids in response to solute concentration gradients [5, 6, 7], is now a well-established active propulsion mechanism, with numerous applications in artificial active matter [8, 9].
@@
The presence of local chemical gradients in microfluidic channels can lead to cooperation or competition with global advective flow, enabling size-dependent colloid transport [10], or particle focussing that can be controlled by manipulating channel geometry and its chemical activity [11]. In particular, the presence of dead-end pores can be used to induce a concentration difference between the main channel and pores large enough to entrain particles [12], or capture and retain them [13], affecting filtration and dispersion in porous media [14]. Active or catalytic pores have also been proposed as local pumps and mixers capable of driving sustained fluid transport without external pressure gradients [15, 16, 17]. When considering catalytic active surfaces, geometric asymmetry alone is sufficient to create heterogeneous concentration fields that can induce propulsion [18, 19] or pumping [20, 21, 22]. Since recent advances in fabrication allow for precise spatial patterning of active regions on surfaces [23, 24], non-uniform coverage can be used together with geometric features to control microscale flow locally.
@@
Here, we focus on a planar wedge-like geometry of dead-end pores, the walls of which are endowed with chemical activity, and which are filled with viscous fluid. Moffatt [25, 26] was the first to examine the effect of this confinement on a flow that emerges in response to a disturbance that acts far away from the tip of the wedge, as well as from a mechanically active sector on the surface, imposing a slip flow on the boundary. In both cases, the celebrated Moffatt eddies emerge as a solution, with an infinite sequence of vortices being created in the fluid, as later seen experimentally by Taneda [27]. Self-similar vortical solutions emerge frequently in externally forced confined flows, such as wedge-shaped trenches with a free surface [28], cone-like geometry [29], electrohydrodynamic flows [30], simulations of driven cavity flows [31, 32], and in ice flows over subglacial mountain valleys [33]. The classical problem of flow actuation by moving boundaries is perhaps best illustrated in the context of corner flows by the Taylor’s scraper problem [34], where one boundary moving boundary ’scrapes’ the fluid, causing its outward motion along the immobile wall. Moffatt’s analysis of corner flow driven by a partial slip on the walls [26], akin to conveyor belts shearing the fluid locally, is an example of (local) mechanical actuation. Here, we explore a similar concept with a chemical actuation mechanism that couples to flow through diffusiophoresis.
@@
The paper is structured as follows. First, in Sec. II, we present the general mathematical framework of diffusiophoretic Stokes flows, describe the geometry of the problem, and the relevant physical quantities. In Sec. III, we discuss the general solutions of the diffusion equation for the solute and the biharmonic equation for the flow stream function in the wedge geometry. We also introduce the formalism of Mellin transforms suited to the geometry considered. Next, in Sec. IV we present analytical solutions for the diffusiophoretic flows induced by uniform coverage of one or both walls with a catalyst. In the following Sec. V we discuss the solute concentration field emerging with a single active sector on one of the walls and the other boundary being absorbing, for different geometric settings. The complementary problem of diffusion with an active sector and a reflective boundary is treated in Sec. VI. Finally, we show in Sec. VII how the solute concentration fields translate to the flow, and obtain the flow field numerically by solving for the stream function. Additionally, in Sec. VIII we discuss the case of the chemical activity of the walls given by analytical functions rather than patches, represented by step-like profiles. We conclude the paper in Sec. IX.
II Phoretic flow generation
@@ II.1 General framework
We adopt a continuum description of phoretic transport, following established theoretical frameworks  [5, 6, 7], to analyse the two-dimensional flow field generated within a wedge formed by two semi-infinite lines starting from one point. The fluid within the wedge is characterised by dynamic viscosity $\eta$, density $\rho$, and contains solute of local concentration $C$, with diffusivity $\kappa$.
The chemical activity $\mathcal{A}$ of a surface $\mathcal{S}$ quantifies the fixed rate of solute release ($\mathcal{A}>0$) or absorption ($\mathcal{A}<0$) on the surface by
$\kappa\bm{n}\cdot\nabla C=-\mathcal{A}\quad\textrm{ on }\mathcal{S}.$ (1)
@@
$\bm{u}=\mathcal{M}(\bm{1}-\bm{n}\bm{n})\cdot\nabla C\quad\textrm{ on }\mathcal{S},$ (2)
@@
$\eta\nabla^{2}\bm{u}$ $=\nabla p,$ (5)
$\nabla\cdot\bm{u}$ $=0.$ (6)
The diffusive Laplace’s problem for the solute concentration $C$ effectively decouples from the hydrodynamic problem and may be solved independently. The concentration $C$ can be used to compute the slip flow along $\mathcal{S}$, eq. (2), which is the boundary condition for the flow field in eq. (5) within the cavity.
Given the boundary conditions that we assume in this problem, and the governing Laplace’s equation, if $C_{0}$ is a constant and the concentration field $C$ is a solution, then $c=C+C_{0}$ is a solution too. The boundary conditions (1), (2) do not depend on the absolute value of the local concentration. For all $C_{0}$, the resulting flow will be the same. This also implies that $c$ can be negative. After shifting by a positive $C_{0}$, we will still obtain a physical solution. In the next sections, we will therefore use the relative concentration $c$.
@@ II.2 Nondimensionalization of the model
Having defined $\mathcal{U}$ and $\mathcal{R}$ as the characteristic velocity and length, respectively, the characteristic concentration is $|\mathcal{A}|\mathcal{R}/\kappa$ and the characteristic pressure is $\eta|\mathcal{AM}|/\mathcal{R}\kappa$. The dimensionless activity and mobility are given by $A=\mathcal{A}/|\mathcal{A}|$ and $M=\mathcal{M}/|\mathcal{M}|$.
The nondimensional governing equations are as follows. For the diffusion of solute, we obtain
@@ III Corner flows
The problem of phoretic flow generation in wedge qualitatively resembles that leading to Moffatt eddies [25, 26]. In the classical problem analysed by Moffatt [25], an infinite sequence of eddies is formed in a wedge-shaped planar domain filled with viscous fluid due to a disturbance acting at asymptotically large distances from the corner. Eddies emerge as a self-similar solution to a second-kind eigenvalue problem for the streamfunction. However, similar flow structures can arise when the fluid is actuated by a moving boundary, as in the famous Taylor’s scraper problem [34], where one of the walls of the wedge-shaped cavity moves with constant speed and drives macroscopic flow. A refined variant of this geometric setting involves motion of a portion of the boundary, when a section of it is endowed with slip velocity, as explored by Moffatt [26]. Using Mellin transform techniques, Moffatt found a solution of this problem for a single moving region (and two regions placed symmetrically), akin to transmission belts mounted within the walls. The solution again involves a sequence of corner eddies. This mechanical example operates using the same principle as one implemented in our case.
Here, we focus on a similar geometry but allow the walls to exhibit chemical activity, which drives the flow. The diffusiophoretic flow generation mechanism leads, in general, to a non-uniform slip velocity profile on the bounding walls. However, for the case of uniform coverage of either one or both walls with catalyst, the induced slip velocity remains constant, therefore directly reducing to the previously obtained results. As we demonstrate analytically in the following, both antisymmetric and symmetric flows can be induced, depending on the specific coverage pattern.
@@ III.1 Solute concentration
We first focus on the problem of diffusion of a solute in the corner domain. Upon introducing cylindrical polar coordinates $(\rho,\theta)$ (see figure 1), the stationary Laplace’s equation for the concentration field $c$ takes the form
@@
if the condition is imposed on the wall $\theta=\alpha$.
Figure 1: Geometry of the diffusiophoretic corner flow setup. An active patch, marked on the $\theta=0$ wall, covering the radial section $\rho\in[a,b]$, releases solute (grey arrows) and generates an inhomogeneous concentration field which drives circulatory flow indicated by schematic streamlines.
@@ III.2 Stokes flow
which becomes the boundary condition for a two-dimensional flow in the fluid domain. In this problem, it is convenient to introduce the Stokes stream function $\Psi(\rho,\theta)$ that obeys a biharmonic equation
@@
$u_{\rho}=\frac{1}{\rho}\frac{\partial\Psi}{\partial\theta}=u_{s},\qquad u_{\theta}=-\frac{\partial\Psi}{\partial\rho}=0,\qquad\text{on {\hbox{\theta= \{0 , \alpha\}}}.}$ (13)
@@ III.3 Solution of Laplace and biharmonic equations in the Mellin space
We shall now introduce the formalism of Mellin transforms to solve the more general case of a single active sector. Indeed, Melling transforms have become the traditional method for solving boundary-value problems in wedge-shaped regions [36, 25, 37]. In cylindrical coordinates, Laplace’s equation takes the form of Eq. (8). Let $\mathds{1}_{\Omega}$ be an indicator function of a set $\Omega$. The Mellin transform $\mathscr{M}_{p}$ of concentration $c$ is denoted as $\bar{c}$
$\mathscr{M}_{p}\{c(\rho,\theta)\}:=\int_{0}^{\infty}\rho^{p-1}c(\rho,\theta)\,d\rho=:\bar{c}(p,\theta).$ (14)
@@
$c(\rho,\theta)=\mathscr{M}^{-1}\{\bar{c}(p,\theta)\}=\frac{1}{2\pi i}\int_{\gamma-i\infty}^{\gamma+i\infty}\bar{c}(p,\theta)\rho^{-p}dp.$ (15)
The condition for the existence of an inverse transform is for $\int_{0}^{\infty}\rho^{\gamma-1}c(\rho,\theta)\,d\rho$ to exist. Unless stated otherwise, $\gamma=0$. Then, for appropriate $c(\rho,\theta)$,
@@
where we introduced an abbreviation for derivatives, $\partial_{\rho}\equiv\partial/\partial\rho$, etc. The Laplace’s equation for concentration, (8), implies that $p^{2}\bar{c}+\partial_{\theta}\bar{c}=0$ and for some $C(p)$, $D(p)$ we have
@@
We now turn to the flow problem. Let $\Psi$ be a stream function. The biharmonic equation, $\nabla^{2}\nabla^{2}\Psi=0$, in cylindrical coordinates and after applying the Mellin transform $\mathscr{M}_{p+4}$ becomes

### v2

# Diffusioosmotic corner flows

@@ Abstract
We study flows generated within a two-dimensional corner by the chemical activity of the confining boundaries. Catalytic reactions at the surfaces induce diffusioosmotic motion of the viscous fluid throughout the domain. The presence of chemically active sectors can give rise to steady eddies reminiscent of classical Moffatt vortices, which are mechanically induced in similar confined geometries. In our approach, an exact analytical solution of the diffusion problem in a wedge geometry is derived and coupled to the diffusioosmotic slip-velocity formulation, yielding the stream function of associated Stokes flow. In selected limiting cases, simple closed-form expressions provide clear physical insight into the underlying mechanisms. Our results open new perspectives for the design of microscale mixing strategies in dead-end pores and cornered microfluidic channels, and offer benchmarks for numerical simulations of confined (diffusio)osmotic systems.
@@ I Introduction
A promising approach exploits phoretic mechanisms, in which surface-generated gradients (of concentration, temperature, etc.) induce effective slip flow on confining surfaces, that in turn gives rise to bulk flow [4]. Diffusioosmosis and diffusiophoresis, the motion of particles and fluids in response to solute concentration gradients [5, 6, 7], is now a well-established active propulsion mechanism, with numerous applications in artificial active matter [8, 9].
The presence of local chemical gradients in microfluidic channels can lead to cooperation or competition with global advective flow, enabling size-dependent colloid transport [10], or particle focusing that can be precisely tuned through the interplay of channel geometry, confinement, and surface chemical activity [11]. Diffusiophoresis has also been used to organize colloids into sharp bands [12], to boost the migration of large particles via imposed solute contrasts [13], and to rectify particle motion to yield motility and pattern formation [14]. Subsequent microfluidic studies revealed steady-state focussing in multicomponent gradients [15]. In colloids-salt mixtures, phoretic effects were shown to affect mixing [16], potentially leading to flow effects ranging from enhanced dispersion to blockage in cellular flows [17]. The coupling of hydrodynamic flows and phoresis can further tune chemotactic and diffusiophoretic spreading [18].
@@
In microscale channels, the presence of dead-end pores can be used to induce a concentration difference between the main channel and pores large enough to entrain particles [19], or capture and retain them [20, 21], affecting filtration and dispersion in porous media [22, 23, 24, 25, 26, 27, 28, 29, 30]. Such diffusiophoretic mechanisms have been shown to enhance solute and particle transport into and out of dead-end pores [31], and to enable size-dependent control of colloidal trapping and release via solute gradients in confined geometries [10], with recent simulations and analyses further quantifying phoretic transport and mixing in narrow channels [32, 33, 34]. Active or catalytic pores have also been proposed as local pumps and mixers capable of driving sustained fluid transport without external pressure gradients [35, 36, 37, 32, 33]. When considering catalytic active surfaces, geometric asymmetry alone is sufficient to create heterogeneous concentration fields that can induce propulsion [38, 39] or pumping [40, 41, 42, 43]. Since recent advances in fabrication allow for precise spatial patterning of active regions on surfaces, for example with catalysts [44, 45], enzymes [46], or surface charges [47, 48], non-uniform coverage can be used together with geometric features to control microscale flow locally.
@@
Here, we focus on a planar wedge-like geometry of dead-end pores, the walls of which are endowed with chemical activity, and which are filled with viscous fluid. Moffatt [49, 50] was the first to examine the effect of this confinement on a flow that emerges in response to a disturbance that acts far away from the tip of the wedge, as well as from a mechanically active sector on the surface, imposing a slip flow on the boundary. In both cases, the celebrated Moffatt eddies emerge as a solution, with an infinite sequence of vortices being created in the fluid, as later seen experimentally by Taneda [51]. Self-similar vortical solutions emerge frequently in externally forced confined flows, such as wedge-shaped trenches with a free surface [52], cone-like geometry [53], electrohydrodynamic flows [54], simulations of driven cavity flows [55, 56], and in ice flows over subglacial mountain valleys [57]. The classical problem of flow actuation by moving boundaries is perhaps best illustrated in the context of corner flows by the Taylor’s scraper problem [58], where one moving boundary ’scrapes’ the fluid, causing its outward motion along the immobile wall. Moffatt’s analysis of corner flow driven by a partial slip on the walls [50], akin to conveyor belts shearing the fluid locally, is an example of (local) mechanical actuation. Here, we explore a similar concept with a chemical actuation mechanism that couples to flow through diffusioosmosis. We note that the problem of flow in the corner geometry can be treated as a limit of linear elasticity theory for a medium enclosed in wedge-like confinement. For the latter, the Green’s functions have recently been found by [59, 60]. For low-Reynolds-number flows, asymptotic behaviour of the Stokeslet singularity in a corner were discussed extensively by [61], and the method of images was used by [62] to explore the dynamics of confined microswimmers.
@@
The paper is structured as follows. First, in Sec. II, we present the general mathematical framework of diffusioosmotic Stokes flows, describe the geometry of the problem, and the relevant physical quantities. In Sec. III, we discuss the general solutions of the diffusion equation for the solute and the biharmonic equation for the flow stream function in the wedge geometry. We also introduce the formalism of Mellin transforms suited to the geometry considered. Next, in Sec. IV we present analytical solutions for the diffusioosmotic flows induced by uniform coverage of one or both walls with a catalyst. In the following Sec. V we discuss the solute concentration field emerging with a single active sector on one of the walls and the other boundary being absorbing, for different geometric settings. The complementary problem of diffusion with an active sector and a reflective boundary is treated in Sec. VI. Finally, we show in Sec. VII how the solute concentration fields translate to the flow, and obtain the flow field numerically by solving for the stream function. Additionally, in Sec. VIII we discuss the case of the chemical activity of the walls given by analytical functions rather than patches, represented by step-like profiles. We conclude the paper in Sec. IX.
II Generation of diffusioosmotic flows
@@ II.1 General framework
We adopt a continuum description of diffusioosmotic transport, following established theoretical frameworks [5, 6, 7], to analyse the two-dimensional flow field generated within a wedge formed by two semi-infinite lines starting from one point. The fluid within the wedge is characterised by dynamic viscosity $\eta$, density $\rho$, and contains solute of local concentration $\tilde{C}$ (number of particles per unit volume), with diffusivity $\kappa$. The chemical activity $\mathcal{A}$ of a surface $\mathcal{S}$ quantifies the fixed rate of solute release ($\mathcal{A}>0$) or absorption ($\mathcal{A}<0$) on the surface by
$\kappa\bm{n}\cdot\nabla\tilde{C}=-\mathcal{A}\quad\textrm{ on }\mathcal{S}.$ (1)
@@
$\tilde{\bm{u}}=\mathcal{M}(\bm{1}-\bm{n}\bm{n})\cdot\nabla\tilde{C}\quad\textrm{ on }\mathcal{S},$ (2)
@@
$\eta\nabla^{2}\tilde{\bm{u}}$ $=\nabla\tilde{P}$, (5)
$\nabla\cdot\tilde{\bm{u}}$ $=0.$ (6)
The diffusive Laplace’s problem for the solute concentration $\tilde{C}$ effectively decouples from the hydrodynamic problem and may be solved independently. The concentration $\tilde{C}$ can be used to compute the slip flow along $\mathcal{S}$ via Eq. (2), which serves as the boundary condition for the flow field in Eq. (5) within the cavity.
We note here that while the fluid is assumed incompressible, in phoretic problems involving the transport of particles, effective particle velocity fields can appear compressible due to concentration-dependent drift. For instance, Raynal et al. [64] showed that diffusiophoretic drift can induce an effectively compressible colloid flow, even in an incompressible solvent, leading to transient particle focusing. Similarly, Chu et al. [65] studied colloid transport under transient solute gradients where spatially varying drift produces apparent compression or expansion, though the underlying fluid remains incompressible.
Given the boundary conditions that we assume in this problem, and the governing Laplace’s equation, if $\tilde{C}_{0}$ is a constant and the concentration field $\tilde{C}$ is a solution, then the excess concentration $\tilde{c}=\tilde{C}-\tilde{C_{0}}$ is a solution too. The boundary conditions (1), (2) do not depend on the absolute value of the local concentration. For all $\tilde{C}_{0}$, the resulting flow will be the same. This also implies that $\tilde{c}$ can be negative. After shifting by a positive $\tilde{C}_{0}$, we thus will still obtain a physical solution. In the next sections, we will therefore replace concentration $\tilde{C}$ with field $\tilde{c}$.
@@ II.2 Nondimensionalization of the model
We have already defined $\mathcal{U}$ and $\mathcal{R}$ as the characteristic velocity and length. The choice of the characteristic length depends on the problem at hand––we discuss this in detail when defining specific problems, but in general it may be set e.g. by the size of an active patch on the wall. Since $\tilde{c}$ is the excess concentration, its variations scale with the magnitude of the gradients produced at confining boundaries, which are induced by surface chemistry. Its natural scale is therefore ${\mathscr{C}=}|\mathcal{A}|\mathcal{R}/\kappa$, where $|\mathcal{A}|$ is the typical magnitude of the chemical surface activity. The characteristic pressure is constructed from the typical velocity scale, given by the mobility and the magnitude of concentration, and reads ${\mathscr{P}=}\eta|\mathcal{AM}|/\mathcal{R}\kappa$. The dimensionless pressure, concentration and velocity fields are thus given by $P={\tilde{P}}/{\mathscr{P}}$, $c={\tilde{c}}/{\mathscr{C}}$, and $\bm{u}={\tilde{\bm{u}}}/{\mathcal{U}}$. Similarly, the dimensionless activity and mobility are given by $A=\mathcal{A}/|\mathcal{A}|$ and $M=\mathcal{M}/|\mathcal{M}|$. Since activity can vary over the surface $\mathcal{S}$, the choice of the characteristic activity is not always obvious. Because $\textrm{Pe},\textrm{Re}\propto|\mathcal{A}|$, choosing the maximal value of $|\mathcal{A}|$ on the surface $\mathcal{S}$ leads to the strongest conditions on Pe and Re. For the diffusion of solute, the nondimensional governing equations are as follows
@@ III Corner flows
The problem of osmotic flow generation in a wedge qualitatively resembles that leading to Moffatt eddies [49, 50]. In the classical problem analysed by Moffatt [49], an infinite sequence of eddies is formed in a wedge-shaped planar domain filled with viscous fluid due to a disturbance acting at asymptotically large distances from the corner. Eddies emerge as a self-similar solution to a second-kind eigenvalue problem for the streamfunction. However, similar flow structures can arise when the fluid is actuated by a moving boundary, as in the famous Taylor’s scraper problem [58], where one of the walls of the wedge-shaped cavity moves with constant speed and drives macroscopic flow. A refined variant of this geometric setting involves motion of a portion of the boundary, when a section of it is endowed with slip velocity, as explored by Moffatt [50]. Using Mellin transform techniques, Moffatt found a solution of this problem for a single moving region (and two regions placed symmetrically), corresponding to transmission belts mounted within the walls. The solution again involves a sequence of corner eddies. This mechanical example operates using the same principle as one implemented in our case.
Here, we focus on a similar geometry but allow the walls to exhibit chemical activity, which drives the flow. The diffusioosmotic flow generation mechanism leads, in general, to a non-uniform slip velocity profile on the bounding walls. However, for the case of uniform coverage of either one or both walls with catalyst, the induced slip velocity remains constant, therefore directly reducing to the previously obtained results. As we demonstrate analytically in the following, both antisymmetric and symmetric flows can be induced, depending on the specific coverage pattern.
@@ III.1 Solute concentration
We first focus on the problem of diffusion of a solute in the corner domain. Upon introducing cylindrical polar coordinates $(\rho,\theta)$ (see figure 1), stationary Laplace’s equation for the concentration field $c$ takes the form
@@
if the condition is imposed on the wall $\theta=\alpha$. For the 2D diffusion equation to be well-posed, we shall require the integrals of both activity functions over the walls to add to zero, as we discuss later in the manuscript.
Figure 1: Geometry of the diffusioosmotic corner flow setup in polar coordinates $(\rho,\theta)$. In a wedge of opening angle $\alpha$, an active patch on the $\theta=0$ wall covering the radial section $\rho\in[a,b]$, releases solute (grey arrows) and generates an inhomogeneous concentration field which drives circulatory flow indicated by schematic streamlines.
@@ III.2 Stokes flow
which becomes the boundary condition for a two-dimensional flow in the fluid domain. In this problem, it is convenient to introduce the two-dimensional stream function $\Psi(\rho,\theta)$ [66, 67] such that in polar coordinates
$u_{\rho}=\frac{1}{\rho}\frac{\partial\Psi}{\partial\theta},\quad u_{\theta}=-\frac{\partial\Psi}{\partial\rho}.$ (12)
Since $\bm{u}=(u_{\rho},u_{\theta})$ satisfies the incompressible Stokes equations (7c)-(7d), $\Psi$ obeys the biharmonic equation
@@
$u_{\rho}=\frac{1}{\rho}\frac{\partial\Psi}{\partial\theta}=u_{s},\qquad u_{\theta}=-\frac{\partial\Psi}{\partial\rho}=0,\qquad\text{for}\ \theta=\{0,\alpha\}.$ (14)
@@ III.3 Solution of Laplace and biharmonic equations in the Mellin space
We shall now introduce the formalism of Mellin transforms [68, 69] to solve the more general case of a single active sector. Indeed, Mellin transforms have become the traditional method for solving boundary-value problems in wedge-shaped regions [70, 49, 50, 71]. In cylindrical coordinates, Laplace’s equation takes the form of Eq. (8). We denote by $\bar{c}(p,\theta)$ the Mellin transform $\mathscr{M}_{p}$ of concentration $c(\rho,\theta)$, where $p$ is the transform variable and
$\mathscr{M}_{p}\{c(\rho,\theta)\}:=\int_{0}^{\infty}\rho^{p-1}c(\rho,\theta)\,\mathrm{d}\rho=:\bar{c}(p,\theta).$ (15)
@@
$c(\rho,\theta)=\mathscr{M}^{-1}\{\bar{c}(p,\theta)\}=\frac{1}{2\pi i}\int_{\gamma-i\infty}^{\gamma+i\infty}\bar{c}(p,\theta)\rho^{-p}\mathrm{d}p.$ (16)
The conditions for the existence of the inverse transform are that for a chosen parameter $\gamma$, the integral $\int_{0}^{\infty}\rho^{\gamma-1}c(\rho,\theta)\,d\rho$ exists and the integration line $(\gamma-i\infty,\gamma+i\infty)$ lies within the strip of analyticity of $\bar{c}(p,\theta)$. When these conditions are satisfied, the inverse is independent of $\gamma$. Unless stated otherwise, we take $\gamma=0$. Then, for appropriate $c(\rho,\theta)$, we have
@@
where we introduced an abbreviation for derivatives, $\partial_{\rho}\equiv\partial/\partial\rho$, etc. The Laplace equation for concentration, (8), implies that $p^{2}\bar{c}+\partial^{2}_{\theta}\bar{c}=0$ and for some $C(p)$, $D(p)$ we have
@@
We now turn to the flow problem. Let $\Psi$ be the stream function. The biharmonic equation, $\nabla^{4}\Psi=0$, in cylindrical coordinates and after applying the Mellin transform $\mathscr{M}_{p+4}$ becomes